OpenStudy (pulsified333):

If f(x)=ax^2+b, f(−1)=−2 and f′(−1)=8, find the coefficients a and b. Answers: a= b=

OpenStudy (pulsified333):

@Michele_Laino

OpenStudy (pulsified333):

@dan815

OpenStudy (michele_laino):

by substitution, we get: \[f\left( { - 1} \right) = a{\left( { - 1} \right)^2} + b = a + b = - 2\]

OpenStudy (pulsified333):

okay

OpenStudy (michele_laino):

now, the first derivative of \(f(x)\), is: \[f'\left( x \right) = 2ax\] please replace \(x=-1\), what do you get?

OpenStudy (pulsified333):

-2a

OpenStudy (pulsified333):

so the coefficient is -2 for a

OpenStudy (michele_laino):

so, we can write: \[ - 2a = 8\]

OpenStudy (michele_laino):

please solve for \(a\)

OpenStudy (pulsified333):

a=-4

OpenStudy (michele_laino):

ok! so, substituting such value for \(a\) into the first equation, we can write: \[\begin{gathered} a + b = - 2 \hfill \\ \hfill \\ - 4 + b = - 2 \hfill \\ \end{gathered} \] please solve for \(b\)

OpenStudy (pulsified333):

2

OpenStudy (michele_laino):

that's right!

OpenStudy (pulsified333):

that makes a lot more sense now :D

OpenStudy (michele_laino):

:)